Phenol, commonly called carbolic acid, is a weak organic acid. If you dissolve 0.892 g of the acid in enough water to make 357 mL of solution, what is the equilibrium hydronium ion concentration? And what is the pH o the solution?
Answer #1

First we must know the concentration of phenol solution.

We have, No. of moles = Mass / Molar mass

\therefore No. of moles of phenol = 0.892 g / 94.11 g/mol = 0.009478 mol

Volume of solution = 357 ml = 0.357 L

We have, Molarity = No. of moles of solute / volume of solution in L

[C6H5OH] = 0.009478 mol / 0.357 L = 0.02655 M

Consider dissociation of phenol in water.

C6H5OH (aq) + H2O \rightleftharpoons C6H5O - (aq) + H3O + (aq)

Equilibrium constant expression for above reaction is K a = [H3O +] [C6H5O - ] / [C6H5OH] = 1.01 \times 10 -10

Let's use ICE table to find out equilibrium concentrations of C6H5O - , H3O + and C6H5OH .

Concentration (M) C6H5OH H3O + C6H5O -
Initial 0.02655
Change -X +X +X
Equilibrium 0.02655 -X X X

Putting above values in Ka formula we get

Ka = (X) (X) / 0.02655 - X = 1.01 \times 10 -10

Phenol is a weak acid, hence we can assume X is very small as compared 0.02655. So we can write 0.02655 -X \simeq 0.02655.

Therefore, K a = X 2 / 0.02655 = 1.01 \times 10 -10

X 2 = 1.01 \times 10 -10 (0.02655)

X 2 = 2.682 \times 10 -12

X = 1.638  \times 10 -06 M = [H3O +] = [C6H5O - ]

We have, pH = -log [H3O +] = - log 1.638  \times 10 -06 = 5.78

ANSWER : Equilibrium hydronium ion concentration = 1.64  \times 10 -06 M

pH of phenol solution = 5.78